OpenStudy (anonymous):

Simplify.square root of 10 times square root of 8

geerky42 (geerky42):

\[\sqrt{10}\sqrt8 = \sqrt{10\cdot8} = \sqrt{(2\cdot5)\cdot(2^3)} = \sqrt{2^4\cdot5}\]

geerky42 (geerky42):

\(=\sqrt{2^4}\sqrt5 = \sqrt{(2^2)^2}\sqrt5 = 2^2\sqrt{5} = 4\sqrt{5}\) See what I did?